is x-2y=10 a function? Explain.
Answer by stanbon(75887)

is x-2y=10 a function?
------------
2y = x-10
y = (1/2)x-5
---
For every value of "x" there is just one value of "y".
So the relation is a function.
====================================
Cheers,
Stan H.
============